# DaCapo: Accelerating Continuous Learning in Autonomous Systems for Video Analytics

### v1

@@ I Introduction
Live video analytics is key to enable seamless functionality of interactive autonomous systems, such as autonomous driving vehicles [8, 30, 20, 72], unmanned aerial vehicles (UAVs) [47, 50, 82], and security surveillance robots [6, 19, 7]. Real-time analysis of live videos is indispensable for comprehending rapidly changing environmental conditions, a pivotal aspect that ensures the successful execution of the autonomous systems’ missions.
@@
To address this issue, continuous learning (CL), an emerging algorithmic methodology, has gained significant attention in recent years [4, 37, 22, 29]. This innovative approach enables autonomous systems to perform inference using lightweight “student” DNN models, which results in reduced resource consumption. However, the limited generality of the lightweight student models gives rise to the data drift problem, where live video data deviates from training data. Continuous learning incorporates constant on-the-spot retrainings, which introduce a new computational challenge, as the systems must now handle not only inference, but also labeling and retraining, as illustrated in Figure 1. We denote these three tasks as continuous learning kernels in this paper.
@@
To evaluate DaCapo’s effectiveness, we use three pairs of (student, teacher) object recognition models (see Table III) and the BDD100K driving video dataset. We model three types of data drift: (1) task change by addition or removal of object labels, (2) transitions between daytime and nighttime, and (3) location change between city and highway. These data drifts are mixed to synthesize size different workload scenarios representing diverse real-world situations that autonomous systems would face. We develop DaCapo using Verilog RTL and synthesize it using Synopsys Design Compiler with TSMC 28nm technology. To cross-validate the cycle-accurate behaviors of DaCapo, we also develop an in-house simulator built upon SCALE-Sim [63]. Our evaluation shows that DaCapo achieves 6.5% and 5.5% higher accuracy than the two state-of-the-art GPU-based continuous learning systems, Ekya and EOMU, respectively. Furthermore, we observe that DaCapo achieves the accuracy improvement, consuming 254$\times$ less energy compared to the GPU baseline. These compelling advantages highlight DaCapo’s role as a bridge that links the algorithmic advancements in continuous learning to their practical deployment on autonomous systems.
@@ II-B Continuously Learning for Video Analytics at Edge
To address the data drift problem, a large body of literature [70, 67, 4, 37] has explored continuous learning, which aims to continuously adapt models to newly appearing data. This approach allows the models to stay up-to-date with the changing data distribution. While continuous learning is a promising approach to reduce compute demand for inference, it introduces new compute load for constant retraining over newly collected data as well as labeling the retraining data by running inference for large, state-of-the-art models, often referred to as teacher models [53]. The challenge arises from the necessity of executing the three distinct yet compute-intensive kernels, each affecting the final accuracy differently, creating a complicated tradeoff space.
@@ III-A Unveiling the Dilemma in Continuous Learning
To understand the performance implications, we perform a preliminary experimental study, using two pairs of Student and Teacher object recognition models: (S:ResNet18, T:WideResNet50) and (S:ResNet34, T:WideResNet101). For the experiments, we use an NVIDIA 3090 and a Jetson Orin’s GPU, representing datacenter-scale and autonomous system-level GPUs respectively. As a baseline for continuous learning systems, we built Ekya under an idealized assumption that the system can optimally utilize the provided GPU resources to obtain the maximum achievable accuracy.
@@ VI-B Hyperparameters for Resource Allocation
Hyperparameters for retraining. For retraining, DaCapo loads a subset of samples from the sample buffer with buffer capacity ${C}_{b}$. The number of samples for retraining (${N}_{t}$) and validation (${N}_{v}$) are predefined, while ${N}_{v}$ is set to half of ${N}_{t}$. These hyperparameters are decided according to the model size, as it has a direct impact on the computational cost required for retraining.
@@
Input: $\hskip 7.5pt{B}_{cur}:\small\textit{Current\enspace sample\enspace buffer}\newline \hskip 6.75pt{W}_{cur}:\small\textit{Current\enspace weight\enspace of\enspace student}\newline \hskip 6.75pt{C}_{b}:\small\textit{Capacity of sample buffer}\newline \hskip 6.75pt{V}_{thr}:\small\textit{Threshold value to detect data drift}\newline \hskip 6.75pt{N}_{t,v}:\small\textit{Retraining hyperparameters}\newline \hskip 6.75pt{N}_{l,ldd}:\small\textit{Labeling hyperparameters}$
@@ 6       ${W}_{cur}\leftarrow\textnormal{{{UpdateWeight}(}}\textnormal{\emph{${W}_{nxt}$}}\textnormal{{)}}$;
7       ${acc}_{val}\leftarrow\textnormal{{{Valid}(}}\textnormal{\emph{${W}_{cur},{D}_{v}$}}\textnormal{{)}}$;
@@ 9       ${O}_{prd}\leftarrow\textnormal{{{GetInferOutputs}(}}\textnormal{\emph{${D}_{l}$}}\textnormal{{)}}$;
10       ${acc}_{lab}\leftarrow\textnormal{{{Evaluate}(}}\textnormal{\emph{${O}_{prd},{D}_{l}$}}\textnormal{{)}}$;
@@ VI-C Spatiotemporal Resource Allocation Algorithm
Algorithm 1 describes the temporal resource allocation algorithm in detail. Line 1-2: DaCapo allocates resources for T-SA and B-SA during the offline phase. Line 4-7: In the retraining phase, DaCapo evaluates the updated model’s weights (${W}_{nxt}$) using a validation dataset (${D}_{v}$) sampled from the current sample buffer (${B}_{cur}$). Note that as validation only involves forward-pass processing on T-SA, its computational cost is significantly lower than that of retraining. The validation accuracy assesses the new model’s adaptability to the accumulated dataset. Line 8-10: During the labeling phase, DaCapo matches inference outputs with corresponding samples to evaluate the current model’s inference accuracy on the newly sampled dataset.
Line 11: If the validation accuracy subtracted from the labeling accuracy is lower than the threshold (${V}_{thr}$), it indicates data drift. Line 12-13: Upon detecting data drift, DaCapo clears the current sample buffer to remove outdated samples, and labeling time is extended to ensure that these better-suited samples are included in the dataset for subsequent phases. Line 14: Once all new samples are labeled, DaCapo updates its sample buffer. With this algorithm, DaCapo adeptly detects and addresses data drift by dynamically allocating labeling time based on the validation accuracy from a segment of the sample buffer.
@@ VII-A Methodology
DNN models.
TABLE III: Specifications of the evaluated deep neural network models for object recognition task.
@@
We evaluate DaCapo’s performance on object classification tasks using six DNN models listed in Table III. These models are paired as student and teacher models: (1) ResNet18 and WideResNet50, (2) ViT-B/32 and ViT-B/16, (3) ResNet34 and WideResNet101.
@@
System emulator implementation. To measure end-to-end accuracy in continuous learning with hardware-based retraining and labeling phases, we develop a system emulator, which uses the accelerator simulator results. It simulates the runtime behaviors of a DaCapo-equipped autonomous system, while the accuracy evaluation is performed on the GPUs. Prior to starting continuous learning, the system generates DNN cycle statistics for the partitioned systolic array, informing the system about the samples processed per phase within a given timeframe. This information guides the execution of continuous learning jobs on the GPU, aligning with the scheduled hardware allocation. By integrating hardware simulation and GPU kernel execution, our approach emulates end-to-end accuracy and the spatiotemporal allocation algorithm within our continuous learning system.
@@
  • RTX3090-CL. This represents the accuracy that can be achieved without any constraints or compression loss by using NVIDIA RTX 3090 GPU and FP32 precision. We use the maximal 30 epochs and 100% sampling rate. This baseline selects the weight parameters that give the maximum accuracy during the retraining epochs.
@@ VII-B Accuracy Comparison
OrinHigh-Ekya and OrinHigh-EOMU. We also compare the Orin GPU baselines with the highest power setting, OrinHigh-Ekya and OrinHigh-EOMU, to examine the full capabilities of NVIDIA’s autonomous system solution. While the two obtain a significant accuracy improvement of respective 11.5% and 13.0% in comparison with OrinLow-Ekya, it is still not closely reaching RTX3090-CL, which suggests potential unleashed benefits of the proposed acceleration solution. Intuitively, when evaluated on the toughest case of (ResNet34, WideResNet101) pair, the two see the largest gains of 45.2% and 43.9% improvements, respectively. We observe a slight difference between Ekya and EMOU since both are based on an assumption that labeling and training are offloaded to remote servers, which makes EOMU’s improved resource allocation method marginally influential on the end accuracy.
@@
On the other hand, when employing DaCapo-Spatiotemporal, the accuracy progresses towards the ideal level (RTX3090-CL), akin to completing the ‘last mile’ in a journey towards optimal performance under the assumption of unlimited computational resources. Concretely, DaCapo-Spatiotemporal achieves significant accuracy gains over OrinHigh-Ekya and OrinHigh-EOMU, which are respectively 6.5% and 5.5% improvements. The results suggest that our hardware-algorithm co-designed acceleration approach effectively unlocks continuous learning’s full capabilities and successfully addresses the “resource wall” challenge.
@@ VII-C In-Depth Analysis on Resource Allocation Algorithm
Accuracy over time. To better analyze the effectiveness of DaCapo-Spatiotemporal over DaCapo-Spatial and the Orin baselines, Figure 10(a) and Figure 10(b) show accuracy change over an extended timeframe, using (ResNet18, WideResNet50) and (ResNet34, WideResNet101) model pairs. In both pairs, DaCapo-Spatiotemporal works better than DaCapo-Spatial and the Orin baselines as they lack ability to promptly adapt to data drift occurrences. The Ekya and Spatial systems use a fixed predetermined window time, which lacks situational awareness within the dataset. However, DaCapo-Spatiotemporal shows better adaptability to environment changes through iterative retraining as required. EOMU is designed to trigger frequent retrainings, showing comparatively better accuracy behaviors than Ekya, while the Orin’s resource insufficiency results in its limited gain.
@@ VII-E Sensitivity Study for Extreme Data Drift Scenarios
Figure 12: Accuracy comparison of DaCapo, EOMU, and Ekya, using extreme data drift scenarios. We use (ResNet18 & WideResNet50) model pair for this experiment.
@@
VII-F Implication of Sampling Algorithms
TABLE V: Averaged accuracy over five scenarios when using two different sampling methods for labeling.
  Model Sampling GMEAN
Method DC-S1 DC-ST2
  ResNet18 & Uniform 81.4% 83.2%
WideResNet50 Active-CL 82.0% (+0.6%) 83.3% (+0.1%)
  ViT-B/32 & Uniform 83.1% 85.5%
ViT-B/16 Active-CL 84.2% (+1.1%) 85.9% (+0.4%)
  ResNet34 & Uniform 78.9% 80.3%
WideResNet101 Active-CL 78.6% (–0.3%) 80.0% (–0.3%)
  • 1
    DaCapo-Spatial
  • 2
    DaCapo-Spatiotemporal
DaCapo uses a uniform sampling method to choose a subset of input data for labeling and retraining. A prior work, Active-CL [81], proposes a novel sampling method customized for continuous learning, which employs the least confidence technique [45] to select important samples to label for retraining. Table V compares the accuracy results of the two sampling methods on DaCapo-Spatial and DaCapo-Spatiotemporal. The results show that Active-CL mostly offers marginal benefits, while even for some cases, producing accuracy loss. Moreover, it is worthwhile to note that sampling method is orthogonal and thus complementary to our proposal.
@@ VIII Related Work
Continuous learning. Continuous learning addresses the critical issue of deployed models becoming outdated over time due to data drift. In the field of machine learning, significant efforts have been directed toward retraining pretrained models with new data without forgetting prior knowledge. The replay-based approach [44, 66, 77, 62, 2, 5] is one such method, integrating valuable historical datasets with new samples for model updates. Alternatively, regularization-based updates [42, 74, 35, 57, 64, 38] augment additional regularization terms during further training to penalize updates on weights critical for previous tasks. However, implementing these solutions often presents practical challenges due to unaddressed training overheads [58, 22]. Distinctively, DaCapo adopts a strategy commonly employed in data streaming scenarios, as exemplified by the video analytics domain.
Live video analytics. Edge devices, positioned near data sources, offer significant benefits, such as bandwidth conservation, reduced latency, and improved privacy by eliminating the need for data transmission to the cloud. However, due to their limited computational capacity, most contemporary edge video analytics systems rely heavily on cloud infrastructure. Prior works [13, 73, 80, 26, 9] aim to use edge devices for preprocessing tasks, such as filtering irrelevant frames or cropping the regions of interest. Several studies [46, 34, 76, 49, 36] propose partitioning DNN computations between edge devices and the cloud. Other works [16, 10, 25, 17] propose enhancing efficiency by caching and reusing previous model inference results. Despite these attempts, the involvement of cloud computing constrains the full potential benefits of edge computing, explaining the increasing development of edge-targeted accelerators in the industry [23, 56, 71, 27, 68, 1]. DaCapo seeks to address these issues. We present a tailored hardware accelerator design that enables autonomous systems to perform independent retraining, thus facilitating a novel continuous learning video analytics system.
@@ IX Conclusion
Most existing continuous learning systems focus solely on training, neglecting the critical yet compute-heavy tasks of inference and labeling, and typically assume the availability of substantial computation resources, which is often impractical on battery-operated, resource-constrained autonomous systems. To tackle this challenge, this work proposes a hardware-algorithm co-designed solution, DaCapo, which enables concurrent executions of retraining, labeling, and inference in a resource-efficient manner. DaCapo achieves this goal by jointly leveraging a spatially-partitionable, precision-flexible accelerator and a spatiotemporal resource allocation algorithm. Our empirical analyses across six real-world scenarios with varying data drifts demonstrate that DaCapo obtains an accuracy improvement of 6.5% over a state-of-the-art system, Ekya, while consuming 254$\times$ less energy. These results confirm that DaCapo is an effective and vital initial step to realize continuous learning on autonomous systems.

### v2

@@ I Introduction
Live video analytics is key to enable seamless functionality of interactive autonomous systems, such as autonomous driving vehicles [38, 10, 93, 42, 35, 25, 87], unmanned aerial vehicles (UAVs) [60, 55, 56, 59, 101], and security surveillance robots [71, 7, 79, 24, 92, 9, 8]. Real-time analysis of live videos is indispensable for comprehending rapidly changing environmental conditions, a pivotal aspect that ensures the successful execution of the autonomous systems’ missions.
@@
To address this issue, continuous learning (CL), an emerging algorithmic methodology, has gained significant attention in recent years [65, 49, 46, 5, 27, 34]. This innovative approach enables autonomous systems to perform inference using lightweight “student” DNN models, which results in reduced resource consumption. However, the limited generality of the lightweight student models gives rise to the data drift problem, where live video data deviates from training data. Continuous learning incorporates constant on-the-spot retrainings, which introduce a new computational challenge, as the systems must now handle not only inference, but also labeling and retraining, as illustrated in Figure 1. We denote these three tasks as continuous learning kernels in this paper.
@@
To evaluate DaCapo’s effectiveness, we use three pairs of (student, teacher) object recognition models (see Table III) and the BDD100K driving video dataset. We model three types of data drift: (1) task change by addition or removal of object labels, (2) transitions between daytime and nighttime, and (3) location change between city and highway. These data drifts are mixed to synthesize different workload scenarios representing diverse real-world situations that autonomous systems would face. We develop DaCapo using Verilog RTL and synthesize it using Synopsys Design Compiler with TSMC 28nm technology. To cross-validate the cycle-accurate behaviors of DaCapo, we also develop an in-house simulator built upon SCALE-Sim [77].
Our evaluation shows that DaCapo achieves 6.5% and 5.5% higher accuracy than the two state-of-the-art GPU-based continuous learning systems, Ekya and EOMU, respectively. Furthermore, we observe that DaCapo achieves the accuracy improvement, consuming 254$\times$ less energy compared to the GPU baseline. These advantages highlight DaCapo’s role in connecting the algorithmic advancements in continuous learning to their practical deployment on autonomous systems. Our tools, including system and hardware simulators and a continuous learning benchmark suite with datasets, are available at https://github.com/casys-kaist/DaCapo.
@@ II-B Continuously Learning for Video Analytics at Edge
To address the data drift problem, a large body of literature [65, 46, 100, 49, 85, 82, 5] has explored continuous learning, which aims to continuously adapt models to newly appearing data. This approach allows the models to stay up-to-date with the changing data distribution. While continuous learning is a promising approach to reduce compute demand for inference, it introduces new compute load for constant retraining over newly collected data as well as labeling the retraining data by running inference for large, state-of-the-art models, often referred to as teacher models [66]. The challenge arises from the necessity of executing the three distinct yet compute-intensive kernels, each affecting the final accuracy differently, creating a complicated tradeoff space.
@@ III-A Unveiling the Dilemma in Continuous Learning
To understand the performance implications, we perform a preliminary experimental study, using two pairs of Student and Teacher object recognition models: (S:ResNet18, T:WideResNet50) and (S:ResNet34, T:WideResNet101). For the experiments, we use an NVIDIA RTX 3090 and an NVIDIA Jetson Orin, representing datacenter-scale and autonomous system-level GPUs respectively. As a baseline for continuous learning systems, we built Ekya under an idealized assumption that the system can optimally utilize the provided GPU resources to obtain the maximum achievable accuracy.
@@ VI-B Hyperparameters for Resource Allocation
Hyperparameters for retraining. For retraining, DaCapo loads a subset of samples from the sample buffer with buffer capacity ${C}_{b}$. The number of samples for retraining (${N}_{t}$) and validation (${N}_{v}$) are predefined, while ${N}_{v}$ is set to one quarter of ${N}_{t}$. These hyperparameters are decided according to the model size, as it has a direct impact on the computational cost required for retraining.
@@
1
Input: $\hskip 7.5pt{B}_{cur}:\textit{Current\enspace sample\enspace buffer}\newline \hskip 7.5pt{W}_{cur}:\textit{Current\enspace weight\enspace of\enspace student}\newline \hskip 7.5pt{C}_{b}:\textit{Capacity of sample buffer}\newline \hskip 7.5pt{V}_{thr}:\textit{Threshold value to detect data drift}\newline \hskip 7.5pt{N}_{t,v}:\textit{Retraining hyperparameters}\newline \hskip 7.5pt{N}_{l,ldd}:\textit{Labeling hyperparameters}$
2
3
@@ 10       ${W}_{cur}\leftarrow\textnormal{{{UpdateWeight}(}}\textnormal{\emph{${W}_{nxt}$}}\textnormal{{)}}$;
11       ${acc}_{v}\leftarrow\textnormal{{{Valid}(}}\textnormal{\emph{${W}_{cur},{D}_{v}$}}\textnormal{{)}}$;
12
@@ 14       ${O}_{prd}\leftarrow\textnormal{{{GetInferOutputs}(}}\textnormal{\emph{${D}_{l}$}}\textnormal{{)}}$;
15       ${acc}_{l}\leftarrow\textnormal{{{Evaluate}(}}\textnormal{\emph{${O}_{prd},{D}_{l}$}}\textnormal{{)}}$;
16
@@ VI-C Spatiotemporal Resource Allocation Algorithm
Algorithm 1 describes the temporal resource allocation algorithm in detail. Line 1-2: DaCapo allocates resources for T-SA and B-SA during the offline phase. Line 4-7: In the retraining phase, DaCapo evaluates the updated model’s weights (${W}_{nxt}$) using a validation dataset (${D}_{v}$) sampled from the current sample buffer (${B}_{cur}$). Note that as validation only involves forward-pass processing on T-SA, its computational cost is significantly lower than that of retraining. The validation accuracy assesses the new model’s adaptability to the accumulated dataset. Line 8-10: During the labeling phase, DaCapo matches inference outputs with corresponding samples to evaluate the current model’s inference accuracy on the newly sampled dataset. Line 11: If the validation accuracy subtracted from the labeling accuracy is lower than the threshold (${V}_{thr}$), it indicates data drift. Line 12-13: Upon detecting data drift, DaCapo clears the current sample buffer to remove outdated samples, and labeling time is extended to ensure that these better-suited samples are included in the dataset for subsequent phases. Line 14: Once all new samples are labeled, DaCapo updates its sample buffer. With this algorithm, DaCapo adeptly detects and addresses data drift by dynamically allocating labeling time based on the validation accuracy from a segment of the sample buffer.
@@ VII-A Methodology
TABLE III: Specifications of the evaluated DNN models.
@@
DNN models. We evaluate DaCapo’s performance on object classification tasks using six DNN models listed in Table III. These models are paired as student and teacher models: (1) ResNet18 and WideResNet50, (2) ViT-B/32 and ViT-B/16, (3) ResNet34 and WideResNet101.
@@
System simulator implementation. To measure end-to-end accuracy in continuous learning with hardware-based retraining and labeling phases, we develop a system simulator, which uses the accelerator simulator results. It simulates the runtime behaviors of a DaCapo-equipped autonomous system, while the accuracy evaluation is performed on the GPUs. Prior to starting continuous learning, the system generates DNN cycle statistics for the partitioned systolic array, informing the system about the samples processed per phase within a given timeframe. This information guides the execution of continuous learning jobs on the GPU, aligning with the scheduled hardware allocation. By integrating hardware simulation and GPU kernel execution, our approach simulates end-to-end accuracy and the spatiotemporal allocation algorithm within our continuous learning system.
@@ VII-B Accuracy Comparison
OrinHigh-Ekya and OrinHigh-EOMU. We also compare the Orin GPU baselines with the highest power setting, OrinHigh-Ekya and OrinHigh-EOMU, to examine the full capabilities of NVIDIA’s autonomous system solution. While the two obtain a significant accuracy improvement of respective 11.5% and 13.0% in comparison with OrinLow-Ekya, it is still not closely reaching DaCapo-Spatiotemporal, which suggests potential unleashed benefits of the proposed acceleration solution. Intuitively, when evaluated on the toughest case of (ResNet34, WideResNet101) pair, the two see the largest gains of 45.2% and 43.9% improvements, respectively. We observe a slight difference between Ekya and EMOU since both are based on an assumption that labeling and training are offloaded to remote servers, which makes EOMU’s improved resource allocation method marginally influential on the end accuracy.
@@
On the other hand, when employing DaCapo-Spatiotemporal, the accuracy progresses towards the highest level among all other continuous learning systems, akin to completing the “last mile” in a journey towards optimal performance under the computational resources of the DaCapo hardware. Concretely, DaCapo-Spatiotemporal achieves significant accuracy gains over OrinHigh-Ekya and OrinHigh-EOMU, which are respectively 6.5% and 5.5% improvements. The results suggest that our hardware-algorithm co-designed acceleration approach effectively unlocks continuous learning’s full capabilities and successfully addresses the “resource wall” challenge.
@@ VII-C In-Depth Analysis on Resource Allocation Algorithm
Accuracy over time. To better analyze the effectiveness of DaCapo-Spatiotemporal over DaCapo-Spatial and the Orin baselines, Figure 10(a) and Figure 10(b) show accuracy change over an extended timeframe, using (ResNet18, WideResNet50) and (ResNet34, WideResNet101) model pairs. In both pairs, DaCapo-Spatiotemporal works better than DaCapo-Spatial and the Orin baselines as they lack ability to promptly adapt to data drift occurrences. The Ekya and DaCapo-Spatial systems use a fixed predetermined window time, which lacks situational awareness within the dataset. However, DaCapo-Spatiotemporal shows better adaptability to environment changes through iterative retraining as required. EOMU is designed to trigger frequent retrainings, showing comparatively better accuracy behaviors than Ekya, while the Orin’s resource insufficiency results in its limited gain.
@@ VII-E Sensitivity Study for Extreme Data Drift Scenarios
Figure 12: Accuracy comparison of DaCapo, EOMU, and Ekya, using two extreme data drift scenarios. We use (ResNet18 & WideResNet50) model pair for this experiment.
@@ VIII Related Work
Continuous learning. Continuous learning addresses the critical issue of deployed models becoming outdated over time due to data drift. In the field of machine learning, significant efforts have been directed toward retraining pretrained models with new data without forgetting prior knowledge. The replay-based approach [81, 96, 76, 2, 6] is one such method, integrating valuable historical datasets with new samples for model updates. Alternatively, regularization-based updates [51, 90, 43, 70, 78, 47] augment additional regularization terms during further training to penalize updates on weights critical for previous tasks. However, implementing these solutions often presents practical challenges due to unaddressed training overheads [72, 27]. Distinctively, DaCapo adopts a strategy commonly employed in data streaming scenarios, as exemplified by the video analytics domain.
Live video analytics. Edge devices, positioned near data sources, offer significant benefits, such as bandwidth conservation, reduced latency, and improved privacy by eliminating the need for data transmission to the cloud. However, due to their limited computational capacity, most contemporary edge video analytics systems rely heavily on cloud infrastructure. Prior works [65, 89, 17, 88, 99, 31, 12] aim to use edge devices for preprocessing tasks, such as filtering irrelevant frames or cropping the regions of interest. Several studies [3, 20, 41, 54, 40, 95, 58, 44] propose partitioning DNN computations between edge devices and the cloud. Other works [100, 21, 13, 30, 22] propose enhancing efficiency by caching and reusing previous inference results. Despite these attempts, the involvement of cloud computing constrains the full potential benefits of edge computing, explaining the increasing development of edge-targeted accelerators in the industry [28, 69, 86, 32, 83, 1]. Furthermore, while there are prior works that propose edge-targeted training solutions [102, 53, 62, 64, 14, 45, 94, 11], they do not take into account the the tradeoff between resource utilization and final accuracy of training and labeling computations. DaCapo addresses these issues by a tailored hardware accelerator design that enables autonomous systems to perform independent retraining, thus facilitating a novel continuous learning video analytics system.
@@ IX Conclusion
Most existing continuous learning systems focus solely on training, neglecting the critical yet compute-heavy tasks of inference and labeling, and typically assume the availability of substantial computation resources, which is often impractical on battery-operated, resource-constrained autonomous systems. To tackle this challenge, this work proposes a hardware-algorithm co-designed solution, DaCapo, which enables concurrent executions of retraining, labeling, and inference in a resource-efficient manner. DaCapo achieves this goal by jointly leveraging a spatially-partitionable, precision-flexible accelerator and a spatiotemporal resource allocation algorithm. Our empirical analyses across six real-world scenarios with varying data drifts demonstrate that DaCapo obtains accuracy improvements of 6.5% and 5.5% over the state-of-the-art systems, Ekya and EOMU, respectively, while consuming 254$\times$ less energy. These results confirm that DaCapo is an effective and vital initial step to realize continuous learning on autonomous systems.
Acknowledgments
We would like to thank Cliff Young, Berkin Akin, and our anonymous reviewers for their review and valuable feedback on the paper. This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) (No.2024-00396013, No.2022-0-01037, No.2018-0-00503) under the Graduate School of Artificial Intelligence Semiconductor (IITP-2024-RS-2023-00256472), Information Technology Research Center (ITRC) support program (IITP-2024-2020-0-01795), and Artificial Intelligence Graduate School Program (KAIST) (No.2019-0-00075), funded by the Korea government (MSIT).